The ages of Hari and Harry are in the ratio 5:7. Four years from now the ratio of their ages will be 3:4. Find their present ages. 
 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the age of Hari be 5x and Harry be 7x.
4 years later,Age of Hari = 5x + 4 Age of Harry = 7x + 4
According to the question,
(5x+4)/(7x+4)=3⁄4
⇒ 4(5x + 4) = 3(7x + 4)
⇒20x+16=21x+12
⇒21x-20x=16-12
⇒x=4Hariage=5x=5×4=20years
Harryage=7x=7×4=28years

Thanks